To: J. Albertine, D. Helms, W. Skinner
cc: Division (M7), FEL Coordination Group
From: F. Dylla
Subject: IRFEL Weekly Report, June 9-13, 1997
Date: June 13, 1997
The second meeting of SURA's Maritime Technical Advisory Committee
was held on June 9-10. The Committee reviewed the laboratory's
FEL development program, the Navy's interest in high energy lasers
and the potential linkage of these interests. Recent assessments
from ONR have significantly impacted the prospects for FY98 Navy
funding. Although, the ONR assessment was positive with respect
to the broad utility of the Jefferson Lab program for industry
and materials science, the Navy has declined near term support
for laser technology for more general defense applications because
of the current tight budgets and higher priority development programs.
MTAC has recommended that we focus all available FY97 Navy funding
and any potential FY98 funding (which will not be determined until
the end of the current Congressional deliberations) towards commissioning
and operating the baseline 1 kW demo and, as funds permit, initiate
the upgrade program to shorter wavelengths and higher powers.
The Navy Program Office reviewed the status of the ITS experiments
and the cryounit warm window development effort during a half
day review following the MTAC meeting. Criteria for declaring
completion of the ITS tests were agreed upon which should allow
start of the ITS cryounit move to the FEL Facility and the ITS
gun move on June 23. (See ITS report for details.)
The preliminary agenda for the June 25-26 LPC workshop was prepared
and distributed to the LPC mailing list. The workshop will focus
on reviewing proposals for initial research applications of the
IR Demo and start-up of the User Labs.
Injector Test Stand
Regarding the photocathode gun:
Most of the week was dedicated to measurements of bunch length
and longitudinal emittance at bunch charges of "zero",
60 pC, and 135 pC. The first step was getting the kicker cavity
to operate in the pulsed mode, and we successfully obtained 300
W/pulse, a value sufficient for the longitudinal- emittance experiments.
We also needed to make a new cathode at mid-week. As of Friday
morning (13 Jun 97) we had obtained the bulk of the data. Though
we had anticipated depleting the cathode during the 135 pC runs
Thursday night, it was still performing well on Friday morning,
and we were therefore planning to augment the 135 pC measurements
over the weekend. We seem to have a bug in the data-reduction
software, in that it gives about a factor of two larger bunch
lengths than the PARMELA simulation code despite otherwise "sensible"
inputs. As a result, we still need to resolve the discrepancy
between the PARMELA algorithm and that of the data-reduction software.
Plans are to finish the 135 pC longitudinal measurements over
the weekend, do a long run next week to monitor dark current versus
time in an effort to gain some additional insight into the field-emission
behavior of the gun, and then begin moving the gun and supporting
systems to the FEL Facility on 23 Jun 97.
The drive laser operated continuously for 97 hours this week in
support of the gun experiments. Its power output was quite stable
over this period. Two brief (less than 1 hour each) intervals
were used to check and optimize the beam alignment. The overall
extinction ratio from the dual set of modulators continues to
be excellent. We did discover that for macropulse lengths longer
than ~10 microseconds, the micropulse intensity is modulated for
~200 microseconds (which, for example, causes corresponding jitter
in the energy spread of the bunches). The problem is due to piezoelectric
resonances in the KD*P crystals used in the modulators. We have
begun discussions with the vendor to determine ways to dampen
Regarding the cryounit:
The cryounit is now warmed up and is being disconnected in preparation
for the 16 Jun 97 move from the ITS to the FEL Facility. Prior
to warm-up we completed testing its waveguides, with forward power
limited to 10 kW. Test results verified that the ceramic warm
windows will support planned commissioning of the IRFEL all the
way through first light with head room available. We also did
a first test of the magnetostrictive tuners on the cryounit.
These tuners are designed for compensating the modest frequency
shifts that will result from electron-beam loading during turn-on
of the injector, which is a relatively slow process. The tuners
have a bandwidth of 0.3 kHz. We found that we could successfully
drive and lock them over their bandwidth, thereby validating their
Regarding development of 50 kW ceramic warm windows:
Production continues with carefully selected ceramics. We expect
to qualify these windows to the 50 kW-level during July. Additional
work continues toward understanding the variability of ceramic
performance. Alternative BeO windows from Northrop Grumman are
in production, with final parts being manufactured by a local
machine shop. Final assembly of these windows is now scheduled
Regarding the beam-transport system:
The injection chicane dipole parts and optical chicane parts are
95% complete. Reverse bend and 180 degree dipole core parts are
in final machining and the first article coils for reverse bends
passed dimensional inspection at the vendor while good progress
was being made on the tooling for the 180 Degree coils. Trial
pole pieces used for prototyping the assembly process are finished
at Everson Electric to begin trials next week.
One and a half sextupole core bodies of the eight required were
cut to the finish dimension by Process Equipment Co. using electro-discharge
machining while the winding mandrel tooling is complete at the
coil vendor. The Trim Quad package was prepared for bidding.
The three, long Panofsky-style quadrupoles (QIs) for the recirculation
dump were continued at New England Electric Co.
Design of the mu metal phasing dipoles was laid out. Broad brush
measurements on the several configurations of a mock-up of an
air core corrector made of four coils in an approximation to a
cosine distribution convinced us that to get the 1% harmonics
or less, we will require a formal cosine distribution dipole.
Magnetic modeling using TOSCA by Steve Lassiter using a wrap-around
end turn configuration was started. Advanced Magnetics Laboratory
(AML) is preparing a preliminary proposal of their own, using
their proprietary winding process, to produce the air core magnets.
Our response will tell them if they are in our technical, price
and delivery comfort zone before they prepare a formal proposal.
The chamber for the 180 degree dipoles was awarded to Master Machine.
Welding of the body on the X Chamber for the Injection Chicane
was completed, the ends machined and we are now checking magnetic
properties in the prototype dipole. The Recirculation Dump's
Y chamber is tacked up. Optical Chicane Chamber plates are now
on hand and being prepared for welding. We are preparing to send
reverse bend chamber parts out for fabrication and are talking
with Stanford Linear Accelerator Center about having them do cleaning,
welding and leak checking.
Design of the final batch of stands and girders for the arcs continued
while a number of girders were installed in the back leg.
Details of the insertable dumps are near sign-off while design
of the beam scraper continued. The straight ahead dump body parts
are done and are being sent out for brazing. Design of the details
of the dump shielding supports continued.
Though much of the instrument air system is installed, we have
finally settled on what compressor to buy and that procurement
Regarding cryomodule fabrication:
Cryomodule assembly continues without issue. The cryounit assembly
area is progressing on the last three cryounits. The cryomodule
assembly area is waiting for the next cryounit, which is scheduled
for delivery next week. The overall schedule leading toward installing
the finished cryomodule in the FEL Facility on 4 Aug 97 remains
tight but doable.
Regarding electron-beam instrumentation and controls (I&C):
Configuring the I&C-related racks in the FEL Facility is an
ongoing activity that is progressing vigorously. As just one
example, the chassis for the beam viewers was installed, and its
documentation (chassis schematic, wiring diagram, front-panel
schematic, art work, fabrication and assembly on both boards)
is now ready for sign-off. Channel assignments are continually
being updated and posted on the www with associated software requirements.
Installation of the safety system for the Facility's clean room
will be completed today (13 Jun 97).
The machine shop had to reinstall one of the quadrupoles on the
wiggler girder after the alignment group noticed that its angle
was wrong. Alignment of the downstream quads and wiggler is complete.
The girder should be installed next week. The vacuum chamber
supports should be finished next week.
We received initial test results from China Lake on the set of
mirrors from Rocky Mountain Instruments (reflectivity ~ 89%).
The mirror figure, surface roughness, and reflectivity are within
their respective specification. However, the ROC appears to be
out of specification. We are consulting with the vendor to check
their metrology. The mirror cans were tested to ensure the input
and output flanges were welded correctly. Our results are in agreement
with the vendor, 8 are better than specification, and 4 are not.
Rather than send the 4 back, we'll use them downstairs where
this specification is not critical. We also tested the mirror-can
camera assembly and found it worked well. The mirror stands are
complete and will arrive on Monday (16 Jun 97), and installation
begins next week. Optical cavity components went through an initial
cleaning. We tested a Cohu charge-coupled-device camera, of the
type we'll use in the accelerator vault, that had been irradiated
with 45 kR of bremsstrahlung from the accelerator. It tested
fine with no dead pixels.
An initial, rudimentary test plan for coherent-synchrotron-radiation
(CSR) experiments was devised and is published on JLab's www site
at ~douglas/FEL/csr_emittance/emittanceindex.html. Sustaining
these experiments early in the commissioning process should not
be a problem, even at 135 pC/bunch because they will be done in
pulsed mode, i.e., at low average current. If the software is
available, we will try to augment the measurements with phase-space
We have made noteworthy progress on self-consistent N-body simulations
that include CSR. The simulations incorporate Gaussian macroparticles,
as opposed to point particles, to eliminate numerical noise.
The key ingredient is a correct calculation of the interaction
force between macroparticles, and we have developed what appears
to be an efficient algorithm for this calculation. To benchmark
the algorithm, we simulated the dynamics of the Gaussian rigid-line
charge using macroparticles and checked to see if the code replicated
our analytical steady-state wakefield, and it rapidly did so with
only 40 macroparticles. Having thereby established the validity
and efficiency of the algorithm in one dimension, we are moving
on to two dimensions. The force is correspondingly more complicated,
but the ingredients of the algorithm remain unchanged. Although
it remains to determine the speed and efficiency of the calculations
in two dimensions, the prognosis for having a self-consistent
simulation tool available in the next few months is good. Of
course, our intent is to use it to predict more accurately emittance
growth in the bends of the transport lattice of the IRFEL and
correlate the results with those of the planned experiments.
We have been following through with our plan for the magnetic-field-measurement
probe for the dipole magnets. The probes are off-the-shelf, obtained
from Rawson-Lush Instrument Co. They have very good absolute
measurement and good reproducibility. However, they do present
a challenge in that the probes are much larger than a normal Hall
probe, and mounting them in the tight magnets will be difficult.
We believe the difficulty is manageable.
Finish work continued at flank speed during this week. The entrance
lobby and stair well was dry walled as well as the entrance double
door area. Ceilings were put on the front and back porch covers.
A large fraction of the earth fill was brought in and bermed
around the facility. Electrical finish work continued. Power
was brought to all the HVAC units and the new clean room. The
elevator crews worked through the week to install doors and controls.
Ceramic tiling continued in the bathrooms; walls are complete
in both men's and women's, flooring in the men's. Showers in
the men's room are 80% complete. Plumbers are preparing to install
fixturing. Touch up painting was done in several areas. Major
painting included break area and stair wells. Preparations are
underway to begin energizing the HVAC units next week. A new
drain was added to the wiggler pit sump. Downstairs Jefferson
Lab crews surveyed the optical cavity tables. Additions were
made to the cable trays in the injector to support cryo and controls.
Dump stands were placed and grouted. The crane hoist was welded
and hung for the cryounit installation. All activities are in
place for moving the cryounit on 6/17. A second rail support
was hung for the gun in preparation for its anticipated move 6/24
(barring unforeseen experimental ITS problems). Cryogenic controls
were pulled and began to be hooked up in preparation for the cooldown.
Leak check continued on the transfer lines. Layouts of the optical
transfer can mounts were passed to the installation crews for
drilling mounting holes on the ceiling. Flooring was laid and
sealed in the clean room. The electric power switch gear was
mounted. Safety system equipment began to be installed in the
clean room. Other safety system equipment was installed all over
the facility: warning lights, ODH sensor wiring, run/safe boxes,
etc., RF installation and hookup of the low level, high voltage
power supply, and high power amplifiers continued. Welding of
the LCW piping continued upstairs. Electrical hookup of the equipment
racks is beginning. Control LAN worked continued.
The crane rails are installed and awaiting load testing. Most of the work on the Clean room is done. The LSS is installed. We will have to wait until June 19 to test the cooling system since chilled water will not be available until then.